OpenStudy (anonymous):

Evaluate the iterated integral. π/4 cosθ ∫ ∫ sinθdrdθ 0 0 Explain please

OpenStudy (unklerhaukus):

are you missing an r from that integrand/

OpenStudy (anonymous):

no thats the entire problem

OpenStudy (unklerhaukus):

\[\text{but the volume element should be } "r\text dr \text d \theta"\]

OpenStudy (anonymous):

This is how the problem is in the textbook I have to take the inner integral part first and then the outside one

OpenStudy (anonymous):

I believe the equation is correct as is. drdtheta gives the order to integrate in, does it not?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

ok

OpenStudy (anonymous):

This should clear up any confusion... http://tutorial.math.lamar.edu/Classes/CalcIII/IteratedIntegrals.aspx

OpenStudy (anonymous):

i looked at that site before but it's not helpful to me

OpenStudy (anonymous):

You just integrate sinθ in regards to dr and as there is no dr don't you treat dr as 1 and the int of 1 = x, then you evaluate on the end points and proceed to take the second integral in regards to theta within the second region.

OpenStudy (unklerhaukus):

\[ \int\limits_0^{π/4} \int\limits_0^{cos \theta} \sin \theta \text{d} r \text{d} \theta\] \[=\int\limits_0^{π/4} r\sin\theta \text d r |_0^{cos\theta} \text d\theta \] \[=\int\limits_0^{π/4}cos{\theta}\sin{\theta}\text d \theta\]

OpenStudy (anonymous):

Yep, that's it ^

OpenStudy (unklerhaukus):

\[=\int\limits_{0}^{π/4}\frac{\sin(2\theta)}{2}\text d {\theta} \]

OpenStudy (anonymous):

the answer is 1/4 but im getting 1/2? i understood everything I was just stuck on where to go from cosxsinx (the antiderivative)

OpenStudy (unklerhaukus):

\[=\frac{1}{4} \cos {(2 \theta)} |_{0}^{π/4} \]

OpenStudy (anonymous):

The derivative of sinx=cosx, so you can use U substitution, thus it should be sin^2(x)/2

OpenStudy (unklerhaukus):

\[=\frac{1}{4}\left(cos(π/2) -\cos(0)\right)\]

OpenStudy (unklerhaukus):

=1/4

OpenStudy (anonymous):

thanks so much i understood it!

OpenStudy (unklerhaukus):

i have used this trig identity \[\sin(2x) = 2 \sin ( x) \cos (x)\] and remember the integral of sin(2x) is cos(2x)/2

OpenStudy (anonymous):

thank you! I'll practice u-sub more and memorizing trig identities.

OpenStudy (unklerhaukus):

i didn't use any substitution

OpenStudy (anonymous):

I'm confused how you went from cosθsinθdθ to sin(2θ)/2dθ. You can use u-sub with cosθsinθ because the derivative of sinx=cosx, so it should be Sinx^2(x)/2, how did you get sin(2θ)/2 ?

OpenStudy (unklerhaukus):

\[\cos(\theta) \sin(\theta)=\sin(2\theta)\] \[\int \sin(\theta) \text d\theta=\cos(\theta)+c\] \[\int \sin(2\theta) \text d\theta=\frac{\cos(\theta)+c}{2}\]

OpenStudy (anonymous):

But, sin(2θ) = 2sin(x)cos(x). You don't have a 2 in front.

OpenStudy (unklerhaukus):

\[\sin(2θ) = 2\sin(x)\cos(x)\]\[\qquad \downarrow\]\[\frac{\sin(2θ)}{2} = \sin(x)\cos(x)\] remember that sine is really \[\cos\theta=\frac{e^{i\theta}+e^{-i\theta} }{2}\] and integrating will not affect the argument of the trigonometric function

OpenStudy (anonymous):

Ah yes. Thanks